# GUARD: Guided Unlearning and Retention via Data Attribution for Large Language Models

### v1

@@ 3.3 Retention-Aware Unlearning Weights
A high retention attribution score for a forget sample $i$ indicates that the sample $(x_{i},y_{i})$ has a positive influence on the performance of the model over the retain set ${\mathcal{D}}_{r}$, or equivalently, its removal leads to a large performance drop on ${\mathcal{D}}_{r}$. We allocate unlearning weights to data samples inversely proportionally to their attribution scores; at the same time, we modulate the variance of the weight distribution by a temperature-controlled softmax. The unlearning weight for a forget sample $i$ equals
@@
where ${a}^{\text{GUARD}}_{i}$ is the GUARD attribution score defined in Eq. 3, $\tau$ is a temperature hyperparameter, and $n_{f}=|I_{f}|$ denotes the size of the forget set. The unlearning weights $\omega^{\text{GUARD}}_{i}$ satisfy the:// 1. Normalization Property: which reads as $\frac{1}{n_{f}}\sum_{i\in I_{f}}\omega^{\text{GUARD}}_{i}=1.$ 2. Inverse Attribution Score Property: The unlearning weights are inversely proportional to the attribution scores of the samples; this mitigates unintended loss of knowledge in ${\mathcal{D}}_{r}$ during unlearning of ${\mathcal{D}}_{f}$. 3. Controlled Variance Property: The degree of weight differentiation is governed by a temperature parameter $\tau$ to enable effective forgetting on ${\mathcal{D}}_{f}$. Comparable forgetting to baselines is guaranteed theoretically (Section 4) and validated empirically (Section 5).
@@ 3.4 Unlearning with GUARD
For a given forget dataset $D_{f}$ and a model parameterized by $\theta$, the novel GUARD unlearning loss objective, $L^{\text{GUARD}}:\mathbb{R}^{p\times q}\to\mathbb{R}$, is defined as
@@
where $\omega_{i}$ is the unlearning weight of datapoint $(x_{i},y_{i})$ from Section 3.3. Following the GA framework in (Eq. 2) as an example, the model update rule under GUARD becomes:
@@
As discussed in Section 3.3, such unlearning weight allocation (Eq. 5) mitigates unintended forgetting of retained knowledge while preserving effective forgetting of undesired information (GUARD objectives for methods other than Gradient Ascent are described in Appendix D.2 and AppendixD.3).
The GUARD algorithm. The descriptions in Sections 3.13.4 are summarized in Algorithm 1, which outlines the implementation of GUARD within the GA unlearning framework (Eq.2). We adopt the standard unlearning protocol with a single forget epoch [51], although the algorithm naturally generalizes to multiple unlearning epochs if needed.
@@
1:Input: Fine-tuned model weights $\theta_{0}$; unlearning rate $\eta$; temperature $\tau$.
@@ 4 Theoretical Guarantees
Let $\psi\in\{0,\text{GA},\text{GUARD}\}$ denote the superscripts corresponding to models trained without unlearning, trained with baseline unlearning (Gradient Ascent, GA), and GUARD, respectively. For a model parameterized by $\theta^{\psi}$, we define the empirical loss over the forget set ${\mathcal{D}}_{f}$ and the retain set ${\mathcal{D}}_{r}$ as:
@@
Let $\overline{g}_{f}:=\frac{1}{n_{f}}\sum_{i\in I_{f}}g_{i}$ and $\overline{g}_{r}:=\frac{1}{n_{r}}\sum_{i\in I_{r}}g_{i}$ denote the average gradients over the forget and retain sets, respectively. We denote the global alignment (defined below) between the two sets as $\kappa:=\langle\overline{g}_{f},\overline{g}_{r}\rangle$ and introduce a normalized alignment metric $\delta_{\kappa}:=\frac{\kappa}{\tau}$, where $\tau$ is the temperature hyperparameter in the GUARD multiplier. Denote the maximum of the norm of model updates upon unlearning as $\delta_{\theta}:=\max_{\psi\in\{\text{GA},\text{GUARD}\}}\|\theta^{\psi}-\theta^{0}\|_{2}$, and define $\delta:=\delta_{\kappa}+\delta_{\theta}$. For each forget sample index $j\in I_{f}$, we have $\kappa_{j}:=\langle\overline{g}_{r},g_{j}\rangle$, whose variance equals
@@
Evaluation metric. We define the Sacrifice Rate $\rho^{\psi}$ to quantify the performance loss on the retain set per unit of performance loss on the forget set:
@@
Assumption 4.1 (Knowledge entanglement between the forget and retain sets).
Assume that the following inequalities hold: 1) $\langle\overline{g}_{r},\overline{g}_{f}\rangle>0$; 2) $\langle\overline{g}_{r},g_{i}\rangle/\tau\ll 1$, $\forall i\in I_{f}$.
Note that gradients can be seen as proxies for information flow [32, 45]. Condition (1) implies a global alignment between the forget and retain sets, implying that their entanglement—unlearning ${\mathcal{D}}_{f}$ typically harms the performance on ${\mathcal{D}}_{r}$. Condition (2) assumes weak alignment between each forget sample and the retain set, enabling selective unlearning without significantly compromising the retained knowledge. The two inequalities jointly yield to $0<\delta_{\kappa}=\kappa/\tau\ll 1$.
Assumption 4.2 (Small updates).
@@
The above assumption is commonly used in influence-function-based data attribution methods [9], and is appropriate whenever unlearning introduces minor perturbations in the model weights.
Assumption 4.3 (Isotropic gradients).
@@
Similar isotropy assumptions were adopted in prior data attribution work for approximating gradient covariance matrices in high-dimensional neural networks [13].
We show next that GUARD significantly reduces the Sacrifice Rate (Theorem 1) by better preserving performance on the retain set (Lemma 4.1) while still achieving effective unlearning on the forget set (Lemma 4.2).
Lemma 4.1 (Retain loss reduction by GUARD).
Under Assumption 4.1-4.3, the loss on the retain set under GUARD is lower than that under GA, i.e.,
$L_{{\mathcal{D}}_{r}}^{\text{GA}}-L_{{\mathcal{D}}_{r}}^{\text{GUARD}}=\frac{\eta}{(1-\delta_{\kappa})\tau}\cdot\sigma^{2}_{\kappa}+\mathcal{O}(\delta^{2}),\text{ where }\eta\text{ is the unlearning rate. }$.
The proof is deferred to Appendix C. Lemma 4.1 shows that the improvement offered by GUARD is proportional to the alignment variance $\sigma^{2}_{\kappa}$ and that it scales with $\frac{\eta}{(1-\delta_{\kappa})\tau}$. A large variance suggests strong interactions between the forget samples and the retain set, which GUARD leverages to reduce the loss. Retention improves with increasing the unlearning rate $\eta$ and decreasing the temperature $\tau$, as long as $\delta_{\kappa}:=\kappa/\tau\ll 1$.
Lemma 4.2 (The forget loss of GUARD is comparable to that of GA).
Under Assumptions 4.1 and 4.2, the difference in forget loss between GUARD and GA is bounded by
@@
The proof is deferred to Appendix C. Lemma 4.2 confirms that GUARD remains effective at unlearning, with a forget loss comparable to that of GA. The gap is small, controlled by $\delta_{\kappa}\ll 1$, and can be further reduced by a smaller squared averaged forget gradient norm $\|\overline{g}_{f}\|_{2}^{2}$ and unlearning rate $\eta$.
@@
Under Assumptions 4.1-4.3, GUARD can be shown to reduce the Sacrifice Rate relative to GA as
@@
The proof is deferred to Appendix C. The reduction is large when 1) the global alignment $\kappa$ is large; 2) the variance $\sigma^{2}_{\kappa}$ is large; 3) the average forget gradient norm is small; and 4) the temperature $\tau$ is small (provided that $\delta_{\kappa}:=\kappa/\tau\ll 1$).
Using Lemma 4.1, Lemma 4.2, and Theorem 1, we provide an in-depth analysis of how GUARD unlearning is influenced by both tunable hyperparameters (the unlearning rate $\eta$ and temperature $\tau$) and intrinsic characteristics of the data/model (knowledge alignment $\kappa$, alignment variance $\sigma_{\kappa}^{2}$, and average forget gradient norm $\|\overline{g}_{f}\|$), in Appendix C.1.
@@ 5 Experiments
5.1 Experimental Setting
The dataset. We evaluate GUARD on the TOFU benchmark [31] for LLM unlearning [15, 14]. TOFU comprises $4,000$ Q-A pairs from $200$ synthetic author profiles. The data is split into: a) Forget/Retain Sets, sampled to unlearn $1\%$, $5\%$, or $10\%$ of the data; b) Generalization Sets (Real-Author, Real-World-Facts) to evaluate downstream retention and generalization.
The backbone LLM. We evaluate the unlearning performance using the two models provided by the TOFU benchmark, Phi-1.5B [23] and Llama-2-7B [43].
Baselines. We adopt the four unlearning baselines provided by TOFU: Gradient Ascent (GA) [42], Gradient Difference (GD) [25], KL Minimization (KM) [7], and Preference Optimization (PO) [29]. Details of these baselines and corresponding GUARD algorithms are provided in Appendix D.1 and D.3.
Evaluation metric. To measure forgetting, retention, and the trade-off between the two, we evaluate the absolute performance and sacrifice rate.
Absolute Performance. Given a dataset $D$ and model $\theta$, we inherit the absolute performance assessment $\epsilon(D;\theta)$ from the TOFU benchmark (e.g., ROUGE-L, Probability, Truth Ratio). A lower $\epsilon({\mathcal{D}}_{f};\theta)$ indicates better forgetting, while a higher $\epsilon({\mathcal{D}}_{r};\theta)$ indicates better retention.
Sacrifice Rate. We introduce the Sacrifice Rate, denoted as $\rho$, defined as
@@ 5.1 Experimental Setting
This metric measures the relative performance degradation on the desired sets per unit of performance decrease on the forget set. A lower $\rho_{\epsilon}$ indicates better relative knowledge retention.
Configuration and hyperparameters. For details, refer to Appendix E.1.
5.2 Results
Following the experimental setup described in Section 5.1, we evaluate the effectiveness of GUARD by primarily focusing on the LLaMA-2-7B backbone (and defer the results for Phi-1.5B to Appendix E.2 as both models exhibit similar behaviours). We report the Sacrifice Rates in Tables 12, and 3, and Absolute Performance metrics in Table 4.
Table 1: Unlearning experiments with LLaMA-7B on 1% of all data points in TOFU. GUARD consistently reduces the Sacrifice Rates across all baselines and evaluation metrics, especially on the Retain Set. It can also lead to negative Sacrifice Rates (i.e., performance gains) on generalization sets, suggesting enhanced generalization via reduction of overfitting on the forget set.
$\rho_{r}$ $\rho_{p}$ $\rho_{t}$ $\rho_{r}$ $\rho_{p}$ $\rho_{t}$ $\rho_{r}$ $\rho_{p}$ $\rho_{t}$
GA 87.87 91.25 289.06 58.15 73.91 172.38 31.54 64.55 119.48
GA + GUARD 49.82 55.32 124.73 39.59 52.36 90.89 17.68 -18.82 -5.93
KM 82.45 79.01 134.99 41.42 18.10 129.15 5.96 19.09 59.17
KM + GUARD 43.42 45.08 24.69 27.16 0.18 55.38 -2.94 -7.08 -1.13
PO 24.59 47.38 94.13 16.39 14.87 61.95 -15.34 5.84 -21.89
PO + GUARD 9.36 31.52 15.99 6.42 4.55 51.36 -17.40 -2.22 -29.21
GD 11.97 12.15 91.65 2.63 4.49 -15.86 -14.29 -17.52 -29.21
GD + GUARD 1.64 8.98 15.51 -1.53 -3.78 -18.04 -17.91 -18.59 -32.02
Table 2: Unlearning experiments with LLaMA-7B on 5% of all data points in TOFU. GUARD consistently reduces the Sacrifice Rates across all evaluation metrics and baselines, with particularly large improvements over baselines that exhibiting weak retention (the most substantial gains are observed against GA, which has the highest Sacrifice Rates).
$\rho_{r}$ $\rho_{p}$ $\rho_{t}$ $\rho_{r}$ $\rho_{p}$ $\rho_{t}$ $\rho_{r}$ $\rho_{p}$ $\rho_{t}$
GA 90.57 118.59 401.01 80.77 88.02 313.25 64.76 37.84 216.97
GA + GUARD 64.31 87.93 237.29 61.73 65.49 236.95 36.65 5.32 165.84
KM 86.15 112.50 337.22 78.74 42.56 180.75 20.50 20.25 98.92
KM + GUARD 60.60 86.79 201.89 58.10 25.17 134.08 -5.24 -11.51 66.66
PO 47.96 104.14 200.59 42.45 25.22 152.90 6.24 -9.37 19.02
PO + GUARD 25.24 82.39 118.83 25.92 13.25 97.87 -4.73 -12.18 -11.29
GD 40.97 79.47 114.25 21.40 -14.44 -17.61 -21.13 -11.53 -11.50
GD + GUARD 18.71 59.30 86.16 10.49 -18.52 -23.77 -27.13 -13.95 -23.74
Table 3: Unlearning experiments with LLaMA-7B on 10% of all data points in TOFU. GUARD offers the most significant improvements under this challenging unlearning scenario when compared to the 1% and 5% forgetting settings. Across all baselines, GUARD markedly mitigates retention degradation, with the most substantial improvement observed on the Retain Set where it reduces $\rho_{t}$ by up to 194.92% relative to GA.
@@ 5.2 Results
Table 4: Impact of GUARD on Absolute Performance evaluated over the Retain and Forget sets. Arrows ($\uparrow$/$\downarrow$) indicate whether higher/lower values are preferred. GUARD consistently improves performance on the Retain Set and maintains comparable performance on the Forget set relative to baseline methods. Hence, GUARD boosts knowledge retention without compromising forgetting.
@@
Effectiveness of GUARD. As shown in Tables 13, GUARD consistently achieves lower sacrifice rates across all datasets, baselines, and unlearning splits, indicating improved retention of unforgotten knowledge. Specifically, GUARD reduces the Sacrifice Rate by up to $194.92\%$ in Truth Ratio on the Retain Set when unlearning 10% of training data. From Table 4, we see that GUARD significantly improves Absolute Performance on the Retain Set while maintaining comparable performance to baselines on the Forget Set, demonstrating its ability to enhance retention without compromising forgetting. Notably, GUARD increases the Truth Ratio by up to $32.08\%$ over the GA baseline on the Retain Set while retaining similar degradation levels on the Forget Set.
Impact of evaluation metrics, baselines, and unlearning percentage. For a fixed unlearning percentage , GA and KM consistently exhibit the highest sacrifice rates (i.e., weakest retention), while GD achieves the lowest. In terms of Absolute Performance (Table 4), GA shows the lowest scores on the Retain Set and the highest on the Forget Set, exhibiting both the worst retention and the most ineffective forgetting. At the same time, GD performs best both with respect to retention and forgetting. Among all evaluation metrics, the Truth Ratio appears to be the most sensitive to unlearning, showing the largest degradation across both relative and absolute evaluations. Furthermore, increasing the proportion of unlearned data consistently leads to higher sacrifice rates, underscoring the growing challenges of retention as unlearning scales increase.
GUARD excels when retention is at its weakest. The improvements achieved by GUARD are most pronounced under challenging conditions. It yields the greatest reductions in Sacrifice Rates and the largest gains in absolute performance for the GA baseline, which suffers from the weakest retention. Similarly, the improvements are most significant for the Truth Ratio, the most sensitive evaluation metric. The largest improvements are observed at the 10% unlearning level.
GUARD preferentially benefits the Retain Set. For a fixed unlearning percentage, baseline, and evaluation metric, GUARD consistently yields the largest improvements over other methods on the Retain Set. This is consistent with it prioritizing the preservation of knowledge in retained data via proxy attribution. Consequently, improvements on the generalization sets (Real World and Real Author Sets) are more indirect. Nonetheless, on these sets we observe cases when the Sacrifice Rates are negative (i.e., have improved performance), suggesting that GUARD not only mitigates degradation but can enhance generalization by alleviating overfitting on the Forget Set.
@@
Additional results. We defer additional experimental findings to the Appendix: 1) Phi-1.5B results (Appendix E.2): We demonstrate that GUARD consistently improves retention across different LLM architectures, including Phi-1.5B. 2) Visualization of proxy attribution and reallocated unlearning power (Appendix E.3): We provide visualizations that confirm the intuition that GUARD reallocates unlearning power based on the proxy attribution scores to mitigate retention losses. 3) Hyperparameter sensitivity analysis (Appendix E.4): We analyze the sensitivity of the temperature parameter $\tau$, supporting our theoretical claim that tuning $\tau$ controls the improvements in retention.

### v2

@@ 3.3 Retention-Aware Unlearning Weights
A large retention attribution score for a Forget Sample $i$ indicates that the sample $(x_{i},y_{i})$ has a strong positive influence on the performance of the model over the Retain Set ${\mathcal{D}}_{r}$ (e.g., its removal leads to a large performance drop on ${\mathcal{D}}_{r}$). We allocate unlearning weights to data samples inversely proportionally to their attribution scores; at the same time, we modulate the variance of the weight distribution by a temperature-controlled softmax. The unlearning weight for a Forget Sample $i$ equals
@@
where as before, ${a}^{\text{GUARD}}_{i}$ is the attribution score defined in Eq. 2, $\tau$ is the temperature hyperparameter, and $n_{f}=|I_{f}|$ denotes the size of the Forget Set. The unlearning weights $\omega^{\text{GUARD}}_{i}$ satisfy the following properties:
1. Normalization Property, which asserts that $\frac{1}{n_{f}}\sum_{i\in I_{f}}\omega^{\text{GUARD}}_{i}=1.$
2. Inverse Attribution Score Property, which asserts that unlearning weights should be inversely proportional to the attribution scores of the samples; this mitigates unintended loss of knowledge in ${\mathcal{D}}_{r}$ during unlearning of ${\mathcal{D}}_{f}$.
3. Controlled Variance Property, which asserts that the degree of weight differentiation is governed by a temperature parameter $\tau$ to enable effective forgetting on ${\mathcal{D}}_{f}$. Comparable forgetting to baselines is guaranteed theoretically (Section 4) and validated empirically (Section 5).
@@ 3.4 Unlearning with GUARD
For a given forget dataset $D_{f}$ and a model parameterized by $\theta$, our novel GUARD unlearning loss objective, $L^{\text{GUARD}}:\mathbb{R}^{p\times q}\to\mathbb{R}$, is defined as
@@
where $\omega_{i}$ is the unlearning weight of datapoint $(x_{i},y_{i})$ from Section 3.3. Using the GA framework in (Eq. 1) as an example, the model update rule under GUARD becomes:
@@
As discussed in Section 3.3, such an unlearning weight allocation (Eq. 4) mitigates unintended forgetting of retained knowledge while it also ensures effective forgetting of undesired information (GUARD objectives for methods other than Gradient Ascent are described in Appendix E.2 and Appendix E.3).
The GUARD algorithm. The descriptions in Sections 3.13.4 are summarized in Algorithm 1, which outlines the implementation of GUARD within the GA unlearning framework (Eq.1). We adopt the standard unlearning protocol with a single forget epoch Yao et al. (2024a), although the algorithm naturally generalizes to multiple unlearning epochs as needed.
@@
1:Input: Initial model weights $\theta_{0}$; unlearning rate $\eta$; temperature $\tau$.
@@ 4 Theoretical Guarantees
Let $\psi\in\{0,\text{GA},\text{GUARD}\}$ denote the superscripts corresponding to models trained without unlearning, trained with standard/baseline unlearning (GA), and GUARD, respectively. For a model parameterized by $\theta^{\psi}$, we define the empirical loss over the Forget Set ${\mathcal{D}}_{f}$ and the Retain Set ${\mathcal{D}}_{r}$ as:
@@
Let $\overline{g}_{f}:=\frac{1}{n_{f}}\sum_{i\in I_{f}}g_{i}$ and $\overline{g}_{r}:=\frac{1}{n_{r}}\sum_{i\in I_{r}}g_{i}$ denote the average gradients over the Forget and Retain Sets, respectively. We denote the global alignment between the two sets as $\kappa:=\langle\overline{g}_{f},\overline{g}_{r}\rangle$ and introduce a normalized alignment metric $\delta_{\kappa}:=\frac{\kappa}{\tau}$, where $\tau$ is the temperature hyperparameter in the GUARD multiplier. Denote the maximum of the norm of model updates upon unlearning as $\delta_{\theta}:=\max_{\psi\in\{\text{GA},\text{GUARD}\}}\|\theta^{\psi}-\theta^{0}\|_{2}$, and define $\delta:=\delta_{\kappa}+\delta_{\theta}$. For each Forget Sample index $j\in I_{f}$, we have $\kappa_{j}:=\langle\overline{g}_{r},g_{j}\rangle$, whose variance equals
@@
Evaluation metric. We define the Sacrifice Rate $\rho^{\psi}$ to quantify the performance loss on the Retain Set per “unit” of performance loss on the Forget Set,
@@
Assumption 1 (Knowledge entanglement between the Forget and Retain Sets).
Assume that the following inequalities hold: (1) $\langle\overline{g}_{r},\overline{g}_{f}\rangle>0$; (2) $\langle\overline{g}_{r},g_{i}\rangle/\tau\ll 1$, $\forall i\in I_{f}$.
Next, note that gradients can be seen as proxies for information flow Nguyen et al. (2025); Wang et al. (2025b). Consequently, condition (1) implies a global alignment between the Forget and Retain Sets, implying that their entanglement—unlearning ${\mathcal{D}}_{f}$ typically harms the performance on ${\mathcal{D}}_{r}$. Condition (2) assumes weak alignment between each Forget Sample and the Retain Set, enabling selective unlearning without significantly compromising the retained knowledge. The two inequalities jointly yield to $0<\delta_{\kappa}=\kappa/\tau\ll 1$.
Assumption 2 (Small updates).
@@
The above assumption is commonly used in influence-function-based data attribution Ellis et al. (2022), and is appropriate whenever unlearning introduces small perturbations in the model weights.
Assumption 3 (Isotropic gradients).
@@
Similar isotropy assumptions were adopted in prior data attribution works for approximating gradient covariance matrices in high-dimensional neural networks Ingram et al. (2023).
We show next that GUARD significantly reduces the Sacrifice Rate (Theorem 3) by preserving performance on the Retain Set (Lemma 1) while effectively unlearning the Forget Set (Lemma 2).
Lemma 1 (Retain loss reduction by GUARD).
Under Assumptions 1-3, the loss on the Retain Set under GUARD is lower than that under GA, i.e.,
$L_{{\mathcal{D}}_{r}}^{\text{GA}}-L_{{\mathcal{D}}_{r}}^{\text{GUARD}}=\frac{\eta}{(1-\delta_{\kappa})\tau}\cdot\sigma^{2}_{\kappa}+\mathcal{O}(\delta^{2}),\text{ where }\eta\text{ denotes the unlearning rate. }$
The proof is deferred to Appendix D. Lemma 1 shows that the improvement offered by GUARD is proportional to the alignment variance $\sigma^{2}_{\kappa}$ and that it scales with $\frac{\eta}{(1-\delta_{\kappa})\tau}$. A large variance suggests strong interactions between the Forget Samples and the Retain Set, which GUARD leverages to reduce the loss. Retention improves with increasing the unlearning rate $\eta$ and decreasing the temperature $\tau$, as long as $\delta_{\kappa}:=\kappa/\tau\ll 1$.
Lemma 2 (The forget loss of GUARD is comparable to that of GA).
Under Assumptions 1 and 2, the difference in forget loss between GUARD and GA is bounded by
@@
The proof is deferred to Appendix D. Lemma 2 confirms that GUARD remains effective at unlearning, with a forget loss comparable to that of GA. The gap is small, controlled by $\delta_{\kappa}\ll 1$, and can be further reduced by a smaller squared averaged forget gradient norm $\|\overline{g}_{f}\|_{2}^{2}$ and unlearning rate $\eta$.
@@
Under Assumptions 1-3, GUARD can be shown to reduce the Sacrifice Rate relative to GA as
@@
The proof is deferred to Appendix D. The reduction is large when 1) the global alignment $\kappa$ is large; 2) the variance $\sigma^{2}_{\kappa}$ is large; 3) the average forget gradient norm is small; and 4) the temperature $\tau$ is small (provided that $\delta_{\kappa}:=\kappa/\tau\ll 1$).
Using Lemma 1, Lemma 2, and Theorem 3, in Appendix D.1 we provide an in-depth analysis of how GUARD unlearning is influenced by both tunable hyperparameters (the unlearning rate $\eta$ and temperature $\tau$) and intrinsic characteristics of the data/model (knowledge alignment $\kappa$, alignment variance $\sigma_{\kappa}^{2}$, and average forget gradient norm $\|\overline{g}_{f}\|$).
@@ 5 Experiments
The dataset. We evaluate GUARD on TOFU Maini et al. (2024b) and MUSE Shi et al. (2024a) benchmarks for LLM unlearning Jia et al. (2024); Ji et al. (2024). TOFU comprises $4,000$ Q-A pairs from $200$ synthetic author profiles. The data is split into: a) Forget/Retain Sets, sampled to unlearn $1\%$, $5\%$, or $10\%$ of the data; b) Generalization Sets (Real-Author, Real-World-Facts) to evaluate downstream retention and generalization. MUSE provides two real-world corpora: a) NEWS (BBC articles) with 0.80.8M/1.61.6M token Forget/Retain Sets; b) BOOKS with Harry Potter books as a 1.11.1M token Forget Set, and FanWiki as 0.50.5M token Retain Set. Each corpus includes verbatim text and knowledge sets with question-answer pairs derived from the original content.
The backbone LLM. Following the original setups of the benchmarks, we evaluate unlearning performance using Phi-1.5B Li et al. (2023) and Llama-2-7B Touvron et al. (2023) on TOFU, Llama-2-7B on MUSE NEWS corpus, and ICLM-7B Shi et al. (2024b) on MUSE BOOKS corpus.
Baselines. We adopt four non-loss reweighting unlearning baselines provided by TOFU and implement GUARD on them. These include: Gradient Ascent (GA) Thudi et al. (2022), Gradient Difference (GD) Liu et al. (2022), KL Minimization (KM) Chundawat et al. (2023), and Preference Optimization (PO) Maini & et al. (2024). Detailed reviews of these baselines and corresponding GUARD algorithms are available in Appendix E.1 and E.3. It is worth noting that since MUSE is a textual dataset and PO is only applicable to QA-Pair datasets, we do not report PO results for MUSE. Additionally, we include results for four recent loss reweighting methods: Negative Preference Optimization (NPO) Zhang et al. (2024a), SimNPO Fan et al. (2025), FLAT Wang et al. (2024b) (we use the TV variant as f-divergence required by FLAT since it achieves the best empirical performance) and SatImp Yang et al. (2025). For fairness of comparisons, we apply GUARD to NPO and compare it against the four loss reweighting methods. We chose NPO as the base method due to its demonstrated stability and excellent performance. However, it is important to note that GUARD can be adapted with equal ease to all other loss reweighting methods to provide further improvements if needed. Details regarding the baselines are provided in Appendix B.
Evaluation metrics. We use tailor-made metrics for TOFU and MUSE benchmarks due to their distinct evaluation criteria.
TOFU Evaluations: To measure forgetting, retention, and the trade-off between the two, we evaluate the Absolute Performance and Sacrifice Rate.
(a) Absolute Performance. Given a dataset $D$ and model $\theta$, we use the Absolute Performance assessment $\epsilon(D;\theta)$ from the TOFU benchmark (e.g., ROUGE-L, Probability, Truth Ratio). A lower value of $\epsilon({\mathcal{D}}_{f};\theta)$ indicates better forgetting, while a higher value of $\epsilon({\mathcal{D}}_{r};\theta)$ indicates better retention.
(b) Sacrifice Rate. We introduce the Sacrifice Rate, denoted as $\rho$, and defined as
@@
This metric measures the relative performance degradation on the desired sets per unit of performance decrease on the Forget Set. The lower $\rho_{\epsilon}$, the better relative knowledge retention.
MUSE Evaluations: Following the original MUSE framework, we evaluate methods according to four key criteria: (a) Forget Set VerbMem. This criteria measures exact text reproduction from the Forget Set. Lower values indicate better unlearning. (b) Forget Set KnowMem. This metric assesses the ability of the model to answer questions about the Forget Set. Lower values indicate better unlearning. (c) PrivLeak. This metric quantifies membership inference vulnerability. Values within [-5%, 5%] indicate successful unlearning, while values outside this range suggest over/under-unlearning. (d) Retain Set KnowMem. This criteria is used to evaluate knowledge retention on the Retain Set. Higher values of the score indicate better utility preservation.
Configuration and hyperparameters. For details, refer to Appendix F.1.
Results. We evaluate the effectiveness of GUARD by focusing on the LLaMA-2-7B backbone, and defer the results for Phi-1.5B to Appendix F.3, as both models exhibit similar behaviors. We report on experimental results for both TOFU (Sacrifice Rates in Table 1 and Absolute Performance metrics in Table 2) and MUSE (four criteria in Table 3). Due to space limitations, in the main text we focus on results for 10% unlearning on TOFU as this is the most challenging scenario, while results for 1% and 5% data unlearning are delegated to Appendix F.2. For the same reason, we only present results for MUSE NEWS and delegate the MUSE BOOKS dataset analyses to Appendix F.4.
Table 1: Unlearning experiments with LLaMA-2-7B on 10% of all data points in TOFU. Across all baselines, GUARD markedly mitigates retention degradation, with the most substantial improvement observed on the Retain Set where it reduces $\rho_{t}$ by up to 194.92% relative to GA.
@@
NPO 85.58 106.24 201.33 42.91 42.47 101.43 8.11 33.24 20.32
SimNPO 76.12 91.22 151.51 36.11 32.94 52.32 5.83 13.51 8.42
FLAT (TV) 73.11 85.46 158.63 29.32 11.27 41.33 5.30 6.26 -12.97
SatImp 71.93 83.11 131.75 31.08 15.94 53.28 8.64 8.85 -5.62
NPO + GUARD 61.12 78.39 118.47 23.88 4.24 11.62 -7.87 -4.55 -13.69
Table 2: Impact of GUARD on Absolute Performance evaluated over the Retain and Forget Sets. Arrows ($\uparrow$$\downarrow$) indicate whether higher/lower values are preferred. GUARD consistently improves performance on the Retain Set and maintains performance on the Forget Set relative to baselines.
@@
Effectiveness of GUARD on TOFU. As shown in Table 1, GUARD consistently offers lower Sacrifice Rates across all datasets and baselines, indicating improved retention of unforgotten knowledge. In particular, GUARD reduces the Sacrifice Rate by up to $194.92\%$ in Truth Ratio on the Retain Set when unlearning 10% of training data. From Table 2, we see that GUARD also significantly improves Absolute Performance on the Retain Set while maintaining comparable performance to baselines on the Forget Set, demonstrating its ability to enhance retention without compromising forgetting. Notably, GUARD increases the Truth Ratio by up to $32.08\%$ over the GA baseline on the Retain Set while maintaining similar degradation levels on the Forget Set as the standard method.
Additionally, GUARD applied on NPO significantly outperforms NPO, SimNPO, FLAT (TV) and SatImp across all three evaluation sets. Specifically, NPO + GUARD achieves substantially lower Sacrifice Rates than SatImp, with reductions of up to $19.18\%$ in Truth Ratio on the Retain Set. Also, NPO + GUARD increases the Truth Ratio by up to 32.08% over SatImp on the Retain Set while retaining similar degradation levels on the Forget Set. This demonstrates that our attribution-driven approach provides superior knowledge retention compared to existing loss reweighting strategies.
Effectiveness of GUARD on MUSE. Based on Table 3, GUARD demonstrates similar effectiveness on MUSE as observed on TOFU. GUARD consistently improves retention performance across all baselines while maintaining effective forgetting, with NPO + GUARD achieving the best overall performance on Retain Set KnowMem ($45.9$) compared to all competing loss reweighting methods. We note that GUARD exhibits somewhat higher PrivLeak scores, which reflects the fact that retention of relevant knowledge naturally makes the model more vulnerable to membership inference attacks. Moreover, this moderate privacy degradation (10.5% increase for NPO + GUARD) is substantially outweighed by the significant utility gains in knowledge retention (a 27.8% improvement on the Retain Set KnowMem) and highly improved forgetting performance (a 22.8% reduction on Forget Set KnowMem), demonstrating that the retention benefits far exceed the privacy loss.
@@
Additional results. 1) Computational efficiency analysis (Appendix F.5): We present a runtime comparisons demonstrating that GUARD introduces minimal computational overhead, with data attribution representing a one-time preprocessing step that amortizes across multiple unlearning operations. 2) Visualization of proxy attribution and reallocated unlearning power (Appendix F.6): We provide visualizations that confirm the intuition that GUARD reallocates unlearning power based on the proxy attribution scores to mitigate retention losses. 3) Hyperparameter sensitivity analysis (Appendix F.7): We analyze the sensitivity of the temperature parameter $\tau,$ supporting our claim that tuning $\tau$ controls the quality of retention.
Table 3: Unlearning experiments with LLaMA-2-7B on MUSE NEWS dataset.
Methods Forget Set VerbMem Forget Set KnowMem PrivLeak Retain Set KnowMem
($\downarrow$) ($\downarrow$) ($\in[-5\%,5\%]$) ($\uparrow$)
NEWS
w/o Unlearn 58.4 63.9 99.8-99.8 55.2
GA 0.0 0.0 5.8 0.0
GA + GUARD 0.0 0.0 6.5 8.4
KM 28.1 50.3 $\mathbf{-95.4}$ 43.5
KM + GUARD 28.5 49.7 101.2-101.2 50.1
GD 4.4 32.4 101.7 27.6
GD + GUARD 2.62 25.9 106.8 45.1
NPO 2.7 56.8 87.7 35.9
SimNPO 2.2 45.1 88.9 38.6
FLAT (TV) 2.3 42.6 81.2 30.3
SimImP 2.5 46.0 92.4 39.5
NPO + GUARD 2.2 43.8 96.9 45.9